A key step in the extraction of iron from its ore is



FeO(s) + CO(g) ⥫⥬ Fe(s) + CO2(g) Kp = 0.403 at 1000°C



This step occurs in the 700°C to 1200°C zone within a blast fur-



nace. What are the equilibrium partial pressures of CO(g) and



CO2(g) when 1.00 atm of CO(g) and excess FeO(s) react in a



sealed container at 1000°C?

1
Expert's answer
2021-11-10T00:29:59-0500

Equilibrium pressure is follows as:

"K_p = \frac{P(CO_2)}{P(CO)} \\

K_p = 0.403"

As the reaction proceeds, it attains equilibrium and hence the equilibrium pressure can be determined using ICE table.



Equilibrium partial pressures of CO(g) and CO2 (g) can be determined substituting values from the ICE table in the equation for Kp

"K_p = \frac{x}{2.05-x}=0.403 \\

x = 0.589"

Equilibrium partial pressure of CO = (2.05 -x) = (2.05 -0.589) = 1.46 atm

Equilibrium partial pressure of CO2 = x = 0.589 atm